# On the interpretation of quantum theory as games between observers and nature played in Minkowski spacetime

### v2

@@ Abstract
In 2019, we introduced games in Minkowski spacetime as a generalization of game theory to special relativity that subsumes games in normal form (spacelike separation) and games in extensive form (timelike separation). Many concepts of the game theory literature, including Nash equilibria and flattened strategic forms, naturally extend to spacetime games. We also emphasized the importance of these games to model quantum experiments, such as Bell experiments and more generally any adaptive measurements, from a decision theory perspective. Subsequent work, named causal contextuality scenarios, suggested formalizing a special case of such games in terms of strategy presheaves. If a natural cover is taken, we show that spacetime games recover causal contextuality scenarios as a special case. Spacetime games provide a visual and intuitive framework that captures the distinction between joint experiments and either-or experiments, so that they are rich enough in their causal structure to imply a natural cover for the corresponding causal contextuality scenario. Based on this observation, we suggest defining the strategy presheaf directly based on the pure strategies (and restrictions thereof) in the reduced strategic form of the spacetime game, and we show that the sheaf property obtains for the games at hand. The argument is simple and similar to event sheaves for the flat case. Finally, we explain how, in the other direction, the failure of the sheaf property on strategy distribution presheaves is consistent with our previous argument that Nash game theory is incompatible with quantum physics. This shows that the insights of spacetime games and causal contextuality scenarios, taken together, can contribute positively to the advancement of the field of quantum foundations.
1 Introduction and background
@@ 1.1 Spacetime games and their extensive form
Fourny (2019a) and Fourny (2020) introduced a new class of games called spacetime games with perfect information, to extend the semantics of game theory to special relativity. In spacetime games, the agents make decisions at various positions in Minkowski spacetime. Spacetime games can be seen as the least common denominator of strategic games on the one hand (spacelike separation), and games in extensive form with perfect information on the other hand (timelike separation).
The same reports explain how these games can be used to model quantum experiments as a game played between observers and nature. An example of a game corresponding to the Bell experiment is shown in Figure 1, where Alice and Bob are the observers, and Alfred, whom we like to visualize as a cat, represents nature.
@@
A canonical injection of the class of spacetime games with perfect information into the class of games in extensive form with imperfect information is also given. Figure 2 shows the game in extensive form with imperfect information that corresponds to the game of Figure 1.
@@ 1.2 Imperfect information and the dashed lines
Imperfect information is visualized with dashed lines, which group nodes into equivalence classes called information sets. The meaning of these dashes lines is that the agent who makes the decision at these nodes does not know at which one of these nodes they are making their decision: they are not fully informed about the past decisions (in the sense of those made at ancestor nodes in the extensive form), which is why it is called imperfect information. Concretely, in Figure 2, Bob is not informed of Alice’s decision ($\{X\}$ or $\{Y\}$), and this is why his two nodes are connected with a dashed line. This has to do with relativity theory: although Alice may have already made her decision, the information has not reached Bob yet, because his decision is spacelike-separated from Alice’s, as can be seen in the spacelike form of Figure 1, which is basically a causal dependency graph in the sense of relativity theory.
The information sets of the extensive form in Figure 2 correspond exactly to those of the spacetime game in Figure 1 (which are singletons, one for each node, because the spacetime game has perfect information). This is always the case for any spacetime games and their extensive form. The generalization of this canonical injection to all spacetime games, even spacetime games with imperfect information, is straightforward and explained in this paper.
@@ 1.3 Imperfect information, locality, and non-contextuality
The semantics of imperfect information and dashed lines has a subtle difference depending on whether one looks at the spacetime game or at the extensive form. In a spacetime game with imperfect information, dashed lines physically represent non-contextuality, connecting identical decisions made in different contexts, whereas in a game in extensive form with imperfect information (in our framework), dashed lines represent either locality or non-contextuality. The spacetime games framework makes explicit that locality is a special case of non-contextuality: in local scenarios, the spacetime game has perfect information and the extensive form has imperfect information, while in more general non-contextualit scenarios, both of them have imperfect information.
@@
Because of the canonical injection of spacetime games into the class of games in extensive form with imperfect information, many well-studied concepts in the game theory literature directly apply to spacetime games. This includes flattening the strategic space with the (reduced) strategic form of a spacetime game (Kuhn, 1950), computing Nash (1951) equilibria, etc.
@@ 1.4 Causal contextuality scenarios
Subsequent work by Abramsky et al. (2024) suggests modelling a subclass of such games, which they call causal contextuality scenarios, as strategy presheaves. As we show in this paper, when natural covers are taken, causal contextuality scenarios are isomorphic to a subclass of spacetime games, which we call alternating spacetime games. This means that spacetime games recover causal contextuality scenarios as a special case, and by transitivity, that they also recover the framework of Gogioso and Pinzani as a special case (Gogioso and Pinzani, 2022)(Gogioso and Pinzani, 2023)(Gogioso and Pinzani, 2023), as well as Hardy models (Hardy, 1993), PR boxes (Popescu and Rohrlich, 1994), and GHZ models (Greenberger et al., 1989).
We also show that a flattened definition of the strategy presheaf based on reduced strategic forms (Kuhn, 1950) emerges naturally and that the sheaf property holds.
@@ 2.1 Definition
To define the mapping, we now give a formal definition of spacetime games. It is equivalent to that given in (Fourny, 2019a) and (Fourny, 2020), but in a more compact form (see also (Baczyk and Fourny, 2023)). In this definition, we do not require perfect information. Imperfect information has the meaning proposed by Kuhn (1950) and extends naturally to spacetime games without affecting the injection into the class of games in extensive form with imperfect information. [^2] This is because merging nodes into information sets in the spacetime game translates directly into merging the corresponding information sets in the extensive form. Spacetime games with perfect information allow for the study of locality, while spacetime games with imperfect information allow for the more general study of non-contextuality.
@@
  • $\mathcal{P}$ is a set of players and $\mathcal{A}$ is a set of actions (in our context, measurement settings and measurement outcomes)
@@
The payoffs associated with the outcomes are part of the semantics of a game but are not the focus of the present paper, which focuses on the game structure. They are also not part of any causal contextuality scenario and are ignored in the bijective mapping.
@@ 2.2 Further examples
Figure 3: A spacetime game with imperfect information that represents a non-contextuality experiment with two settings to pick out of three. Bob has one information set with three actions (the three contexts XY, XZ, and YZ). Alfred has 3 information sets X, Y, and Z for each measurement, with two actions each, 0 and 1. This is also known in the literature as a cyclic system of rank 3. Imperfect information in spacetime games captures a more general of non-contextuality than the special case of locality.
@@
Figure 4: The game in extensive form with imperfect information that corresponds to the spacetime game shown in Figure 3. It has exactly the same information sets: XY, YZ, and XZ for Bob, and X, Y, and Z for Alfred.
Figure 5 shows an example of a spacetime game with perfect information that has a causal bridge and the corresponding extensive form is shown in Figure 6. Spacetime games can be much more complex and mix spacelike-separation as in Figure 1, timelike-separation as in Figure 5, and imperfect information for explicit contexts as shown in Figure 3.
@@
Figure 6: The game in extensive form with, in this case, perfect information, that corresponds to the spacetime game shown in Figure 5. It happens to have the same tree shape because there is no spacelike separation.
@@
Figure 8: The game in extensive form with imperfect information that corresponds to the spacetime game shown in Figure 7. It has exactly the same information sets: A for Bob, and X, Y, W, and Z for Alfred.
@@ 2.4 Histories
We now give a compact definition of the histories of a spacetime game:
@@
Fourny (2020) uses the term “contingency coordinates” of a node $N$ to denote the minimal history not containing $N$ that coincides with any history in which $N$ is assigned an action (i.e., not $\bot$). In this paper, we use the term “causal bridge” (Fourny et al., 2014) [^3] The idea of a decision happening only if a certain causal past is secured, and otherwise not, can be traced back to (Dupuy, 1992) and (Dupuy, 2000), and we retain his terminology “causal bridge”. of a node $N$ to specifically designate the predecessors of $N$ and the labels on the edges from the predecessors to $N$. The two differ in that the contingency coordinates involve the entire ancestry, while the causal bridge only involves (directly connected, one level up) predecessors.
@@
We always assume that spacetime games are valid, as we consider validity to be part of their definition. Given a complete history (outcome) $z$, we write
@@
Without loss of expressive power, we assume that games do not contain unused information sets, i.e., there is no information set that never gets activated; otherwise, unused information sets can simply be pruned from the game. [^4] The same assumption will be made for causal contextuality scenarios: No measurement never gets activated, otherwise such measurements can be pruned from the causal contextuality scenario. Likewise, we assume there are no unused actions in $\mathcal{A}$.
@@ 3 Results
Thus, spacetime games and causal contextuality scenarios stand in bijection for a broad class of experimental protocols, which can be either characterized as alternating spacetime games or as causal contextuality scenarios with “good covers” (see open question 7.3 by Abramsky et al. (2024)), whereby we suggest that a cover is “good” is it corresponds to the natural cover implied by a spacetime game structure [^5] Abramsky et al. (2024) more recently referred to such good covers as “causally secured covers.”.
@@ 3.1 Alternating spacetime games
We now give a characterization of a subclass of spacetime games called “alternating spacetime games”, for which we will show that there is a mapping with the corresponding causal contextuality scenarios.
@@
We include, in the definition of the alternating property, the interpretation of the assumed features of the game in terms of quantum-experimental protocols, to facilitate the understanding of the mapping that will be given further down.
@@ 3.3 Standardization of Bob’s actions in alternating spacetime games
Without loss of generality, we can put constraints on Bob’s actions in an alternating spacetime game to facilitate their interpretation as contexts, like so:
@@ 3.4 Causal contextuality scenarios
We recall the definition of a causal contextuality scenario as given by Abramsky et al. (2024):
@@
A causal contextuality scenario $\Gamma$ is a quadruplet $(X,O,\vdash,\mathcal{C})$ where X is a set of measurement settings, $O=(O_{x})_{x\in X}$ is a family of sets of measurement outcomes, indexed by settings, and $\vdash$ is an enabling relation between a section on a subset of $X$ (a “consistent set of events”) and settings, and $\mathcal{C}$ is a family of subsets of $X$ whose union is X.
@@
Furthermore, we restrict our attention in this paper to covers that satisfy the antichain property. It is, however, possible to extend the mapping naturally even if the antichain property does not hold (e.g., by allowing unused actions at Bob’s nodes).

### v3

# Spacetime games subsume causal contextuality scenarios

@@ Abstract
We show that causal contextuality scenarios are equivalent to a subset of the formerly published spacetime games, which generalize game theory to decisions arbitrarily located in Minkowski spacetime. This insight leads to certain constructs and proofs being shorter, simpler, and more intuitive when expressed in the spacetime game framework than in the causal contextuality scenario framework. This shows that the insights of both frameworks taken together can contribute positively to advancing the field of quantum foundations.
keywords:
Game theory, Quantum theory, Contextuality, Locality, Bell inequalities, Special relativity
1 Introduction
Quantum theory has been challenging the physics community for a century now. The question remains open of whether there exists an extension theory, as envisioned by Einstein, Podolsky, and Rosen (EPR) [1], which is able not only to predict the same statistical distributions as the Born rule for repeated experiments—they are observed in Nature with very high precision—but also to predict each individual measurement outcomes.
But the way such a theory can look like is constrained. Indeed, two assumptions [2] are made for Bell inequalities to hold [^1] Determinism is derived from these two with a short argument made by Bell [2], thus the companion adjective “local-realist.”:
  • Einsteinian locality: no element of information propagates faster than light.
  • Measurement independence: the choice of a measurement setting, by the experimenter, is exogeneous and, as such, independent of anything not in its future light cone.
Since there is experimental evidence [3] that Bell inequalities are broken by Nature, it follows that one of these two assumptions does not hold. Most of the quantum physics literature so far has focused on assuming measurement independence and dropping Einsteinian locality. Little, but not nothing, has been published on the second possibility, namely, that measurement independence does not hold.
Perhaps this is related to the fact that this assumption of measurement independence is also commonly called “free choice” or “free will” and can lead to strong or emotional philosophical opinions on the matter: we are scientists, but we are also human beings. Yet, it is so that measurement independence is an assumption that can be formalized mathematically in terms of the statistical independence and/or the counterfactual independence (the two are distinct but related in subtle ways [4]) of random variables, and the scientific approach dictates that we must consider the possibility that measurement independence does not hold with as much reason as the possibility that it does. The risk of failing to do so is that an extension theory exists but remains undiscovered because we do not commit any resources to finding it.
One of the most interesting developments in the quantum foundations is the realization that many surprising and disturbing aspects of quantum theory can be abstracted away from physical experiments, and defined purely in terms of which experiments a physicist decides to carry out, and which outcomes they observe. Furthermore, one does not need infinite lists of choices: EPR setups are already interesting with two agents, two settings per agent, and two outcomes per agent and setting (2-2-2). We demonstrated [5] that one can model and view an experimental setup as a game between observers and the universe, Nature being an economic agent that minimizes action (principle of least action). Baczyk and Fourny [6] have given a qualitative argument that the Nash equilibrium [7] paradigm, which has been the mainstream framework of game theory for 75 years, corresponds to the conditions that lead to Bell inequalities:
  • locality corresponds to imperfect information (dashed lines);
  • realism corresponds to the strategic form that can be built from a game in extensive form;
  • measurement independence corresponds to unilateral deviations;
  • a local hidden-variable model corresponds to a mixed strategy (which is a statistical mixture of pure strategies) in the strategic form.
With this insight, one learns that any extension theory of quantum physics, with this game interpretation, cannot follow the Nash paradigm, and must follow instead a non-Nashian game theory paradigm, such as the Perfect Prediction Equilibrium or the Perfectly Transparent Equilibrium [4].
Subsequent work by Abramsky et al [8] builds on our idea that quantum experimental setups can be seen as games between observers and the universe and models quantum experiments as causal contextuality scenarios. In this paper, we show that our framework subsumes causal contextuality scenarios, i.e., that causal contextuality scenarios are equivalent to a subclass of spacetime games. We provide the formal mapping in both directions. We conclude by showing that the bridge between the two formalisms is mutually beneficial and leads to a clearer understanding of a decision-theoretical view of quantum physics. In particular, we discuss how certain seemingly complex constructs, reasonings, or proofs involving causal contextuality scenarios become one-line arguments in the spacetime game framework.
2 Background
But before we do so, we are first going through the spacetime game framework with a few examples.
@@ 2.1 Spacetime games and their extensive form
Game theory has two main game structures [9]: the normal form and the extensive form. As explained by Baczyk and Fourny [10][6], the normal form corresponds to spacelike-separated decisions, while the extensive form corresponds to timelike-separated decisions and even adaptive measurements. A general framework combining both, i.e., for decisions with arbitrary positions in spacetime, did not exist before 2019 to the best of our knowledge. If we are to consider generic quantum experimental setups with labs running local experiments placed at arbitrary positions in spacetime, and sending quantum states to one another—this corresponds to a fixed causal order in the process matrix framework [11]—, we need an extension and unification of the normal form and extensive form.
Thus, a new class of games called spacetime games with perfect information was introduced [5] [10]. It extends the semantics of game theory to special relativity. In spacetime games, the agents make decisions at various positions in Minkowski spacetime. Spacetime games can be seen as the least common denominator of games in normal form on the one hand (spacelike separation), and games in extensive form with perfect information on the other hand (timelike separation).
@@
The same documents explain how these games can be used to model quantum experiments as a game played between observers and nature. An example of a game corresponding to the Bell experiment is shown in Figure 1, where Alice and Bob are the observers, and Alfred, whom we like to visualize as a cat, represents nature.
A canonical injection of the class of spacetime games with perfect information into the class of games in extensive form with imperfect information is also given. Figure 2 shows the game in extensive form with imperfect information that corresponds to the game of Figure 1. The two have equivalent semantics.
This is important because it means that spacetime games, when converted to games in extensive form with imperfect information, inherit several decades’ worth of research, constructs, and proofs carried out on games in extensive form with imperfect information for the past 75 years. This is key to our claim that several constructs and proofs behind causal contextuality scenarios can be significantly simplified. In other words, spacetime games provide an intuitive visual representation for quantum experimental setups but also a conversion algorithm to represent such setups as games in extensive form with imperfect information.
@@ 2.2 Imperfect information and the dashed lines
Imperfect information is visualized with dashed lines, which group nodes into equivalence classes called information sets. The meaning of these dashed lines is that the agent who makes the decision at these nodes does not know at which one of these nodes they are making their decision: they are not fully informed about the decisions made at ancestor nodes [^2] These ancestor nodes need not all be in the past, they can also be spacelike-separated in the spacetime game. Since information cannot propagate faster than light, it is not possible to be informed about decisions made at spacelike-separated locations. in the extensive form. This is why it is called imperfect information. Concretely, in Figure 2, Bob is not informed of Alice’s decision ($\{X\}$ or $\{Y\}$), and this is why his two nodes are connected with a dashed line. This has to do with relativity theory: although Alice may have already made her decision, the information has not reached Bob yet because his decision is spacelike-separated from Alice’s, as can be seen in the spacelike form of Figure 1, which is a causal dependency graph in the sense of relativity theory.
The information sets of the extensive form in Figure 2 correspond exactly to those of the spacetime game in Figure 1 (which are singletons, one for each node because the spacetime game has perfect information). This is always the case for any spacetime games and their extensive form. The generalization of this canonical injection to all spacetime games, even spacetime games with imperfect information, is straightforward and explained in this paper.
@@ 2.3 Imperfect information, locality, and non-contextuality
The semantics of imperfect information and dashed lines have a subtle difference depending on whether one looks at the spacetime game or its extensive form. In a spacetime game with imperfect information, dashed lines physically represent non-contextuality, connecting identical decisions made in different contexts, whereas in a game in extensive form with imperfect information (at least those built from a spacetime game), dashed lines represent either locality or non-contextuality. The spacetime game framework makes explicit that locality is a special case of non-contextuality: in local scenarios, the spacetime game has perfect information and the extensive form has imperfect information, while in more general non-contextuality scenarios, both of them have imperfect information.
@@
Because of the canonical injection of spacetime games into the class of games in extensive form with imperfect information, many well-studied concepts in the game theory literature apply to spacetime games. This includes flattening the strategic space with the (reduced) strategic form of a spacetime game [10][12], computing Nash equilibria [7], mixed strategies (which we will see are also global sections of empirical models), etc.
@@ 2.4 Causal contextuality scenarios
Subsequent work by Abramsky et al [8] suggests modeling a subclass of such games, which they call causal contextuality scenarios, as strategy presheaves. As we show in this paper, when natural covers are taken, causal contextuality scenarios are isomorphic to a subclass of spacetime games, which we call alternating spacetime games. This means that spacetime games recover causal contextuality scenarios as a special case, and by transitivity (based on the claims made in [8]), that they also recover the framework of Gogioso and Pinzani as a special case [13], as well as Hardy models [14], PR boxes [15], and GHZ models [16].
We also show that a flattened definition of the strategy presheaf based on reduced strategic forms [12] emerges naturally and that the sheaf property holds, and that temporal contextuality scenarios are subsumed by games in extensive form with perfect information.
@@ 3.1 Definition
To define the mapping, we now give a formal definition of spacetime games. It is equivalent to that previously published [5] [10], but in a more compact form (see also [6]). In this definition, we do not require perfect information. Imperfect information has the meaning proposed by Kuhn [12] and extends naturally to spacetime games without affecting the injection into the class of games in extensive form with imperfect information. [^3] This is because merging nodes into information sets in the spacetime game translates into merging the corresponding information sets in the extensive form. Spacetime games with perfect information allow for the study of locality, while spacetime games with imperfect information allow for the more general study of non-contextuality.
@@
  • $\mathcal{P}$ is a set of players and $\mathcal{A}$ is a set of actions (in our context, actions are measurement settings and measurement outcomes)
@@
The payoffs associated with the outcomes are part of a game’s semantics but are not the focus of the present paper, which focuses on the game structure. They are also not part of any causal contextuality scenario and are ignored in the bijective mapping.
@@ 3.2 Further examples
Figure 3: A spacetime game with imperfect information that represents a non-contextuality experiment with two settings to pick out of three. Bob has one information set with three actions (the three contexts XY, XZ, and YZ). Alfred has 3 information sets X, Y, and Z for each measurement, with two actions each, 0 and 1. This is also known in the literature as a cyclic system of rank 3. Imperfect information in spacetime games captures the more general case of non-contextuality than the special case of locality.
@@
Figure 4: The game in extensive form with imperfect information that corresponds to the spacetime game shown in Figure 3. It has the same information sets: XY, YZ, and XZ for Bob, and X, Y, and Z for Alfred.
Figure 5 shows an example of a spacetime game with perfect information that has a causal bridge (adaptive measurement) and the corresponding extensive form is shown in Figure 6. Spacetime games can be much more complex and mix spacelike-separation as in Figure 1, timelike-separation as in Figure 5, and imperfect information for explicit contexts as shown in Figure 3.
@@
Figure 6: The game in extensive form with, in this case, perfect information, that corresponds to the spacetime game shown in Figure 5. It happens to have the same tree shape because there is no spacelike separation. Such games correspond to temporal contextuality scenarios.
@@
Figure 8: The game in extensive form with imperfect information that corresponds to the spacetime game shown in Figure 7. It has the same information sets: A for Bob, and X, Y, W, and Z for Alfred.
@@ 3.4 Histories
We now give a compact definition of the histories of a spacetime game (which is a compact form of that given in [10][6]):
@@
In [10], we use the term “contingency coordinates” of a node $N$ to denote the minimal history not containing $N$ that coincides with any history in which $N$ is assigned an action (i.e., not $\bot$). In this paper, we use the term “causal bridge” [17] [^4] The idea of a decision happening only if a certain causal past is secured, and otherwise not, can be traced back to Dupuy [18][19], and we retain his terminology “causal bridge”. of a node $N$ to specifically designate the predecessors of $N$ and the labels on the edges from the predecessors to $N$. The two differ in that the contingency coordinates involve the entire ancestry, while the causal bridge only involves (directly connected, one level up) predecessors.
@@
We always assume that spacetime games are valid, as we consider validity to be part of their definition. Given a complete history (outcome) $z$, we write:
@@
Without loss of expressive power, we assume that games do not contain unused information sets, i.e., there is no information set that never gets activated; otherwise, unused information sets can be pruned from the game. [^5] The same assumption will be made for causal contextuality scenarios: There is no measurement that never gets activated, otherwise such measurements can be pruned from the causal contextuality scenario. Likewise, we assume there are no unused actions in $\mathcal{A}$.
@@ 4 Results
Thus, spacetime games and causal contextuality scenarios stand in bijection for a broad class of experimental protocols, which can be either characterized as alternating spacetime games or as causal contextuality scenarios with “good covers” (see open question 7.3 by Abramsky et al [8]), whereby we suggest that a cover is “good” is it corresponds to the natural cover implied by a spacetime game structure.
@@ 4.1 Alternating spacetime games
Spacetime games, but also games in normal or extensive form with or without perfect information, are very general and cover many different scenarios, from microeconomics to leisure (Chess, etc). Here, we are looking at specific kinds of spacetime games in which experimenters and the universe alternate in choosing settings and outcomes. We thus now give a characterization of a particular relevant subclass of spacetime games called “alternating spacetime games”, for which we will show that there is a mapping with the corresponding causal contextuality scenarios.
@@
In the definition of the alternating property, we interpret the game’s assumed features in terms of quantum-experimental protocols, to facilitate the understanding of the mapping that will be given further down.
@@ 4.3 Standardization of Bob’s actions in alternating spacetime games
Without loss of generality, we can put constraints on the names given to Bob’s actions in an alternating spacetime game to facilitate their interpretation as contexts, like so:
@@
An alternating game A causal contextuality scenario
An alternating game that is in normal form A flat contextuality scenario
An alternating game whose extensive form has perfect information A temporal contextuality scenario
The reduced strategic form of the spacetime game Flattening construction
A pure strategy of the reduced strategic form A global strategy
A mixed strategy of the reduced strategic form The global section of an empirical model (if it exists)
(Dupuysian) causal bridge Causally secured
The projection of a complete history to Bob A “good’) cover
@@ 4.4 Causal contextuality scenarios
We recall the definition of a causal contextuality scenario as given by Abramsky et al [8]:
@@
A causal contextuality scenario $\Gamma$ is a quadruplet $(X,O,\vdash,\mathcal{C})$ where X is a set of measurement settings, $O=(O_{x})_{x\in X}$ is a family of sets of measurement outcomes, indexed by settings, $\vdash$ is an enabling relation between a section on a subset of $X$ (a “consistent set of events”) and settings, and $\mathcal{C}$ is a family of subsets of $X$ whose union is X.
@@
Furthermore, this paper restricts its attention to covers that satisfy the antichain property. It is, however, possible to extend the mapping naturally even if the antichain property does not hold (e.g., by allowing unused actions at Bob’s nodes).